# SPHINX: Structural Prediction using Hypergraph Inference Network

### v1

@@ 3.2 Discrete constrained sampling
The hypergraph predictor module, as described above, produces a probabilistic incidence matrix $\mathbf{P}\in\mathbb{R}^{N\times M}$, where each element $p_{i,j}$ denotes the probability of a node $i$ being part of the hyperedge $j$. However, the standard hypergraph neural network architectures are designed to work with discrete rather than probabilistic structures.
Previous work [43] employs Gumbel-Softmax [10] to sample from a categorical distribution in a differentiable way. However, these techniques sample each element in the hyperedge independently, without any control on the cardinality of the hyperedge. This leads to unstable optimisation, that requires additional training strategies such as specific sparsity regularisation.
To address this issue, we are leveraging the recent advancement in constrained $k$-subset sampling [8, 46, 9]. These methods were successfully used to tackle discrete problems such as combinatorical optimisation, learning explanations and, more recently, rewiring graph topology [47]. Different than classical differentiable samplers [10], the $k$-subset sampler would produce a subset of size exactly $k$, equipped with a gradient estimator useful for backpropagation.
@@ 3.3 Hypergraph processing
Producing a discrete hypergraph structure and being able to propagate the gradient through the entire pipeline, enable us to process the resulting latent hypergraph with any existing architecture designed for higher-order representations. In the recent years, several architectures were developed for hypergraph-structured input [26, 48, 27, 28]. Most of them are following the general two-stage message-passing framework. In the first stage, the information is sent from nodes to the hyperedges using a permutation-invariant operator $z_{j}=f_{V\rightarrow E}(\{x_{i}|v_{i}\in e_{j}\})$. On the second stage the messages are sent back from hyperedge to nodes $x_{i}=f_{E\rightarrow V}(\{z_{j}|v_{i}\in e_{j}\})$.
In our experiments we use a similar setup to the one proposed in [28], in which the two functions $f_{V\rightarrow E}$ and $f_{E\rightarrow V}$ are implemented as DeepSets [29] models.
$f_{V\rightarrow E}(S)=f_{E\rightarrow V}(S)=\text{MLP}(\sum_{s\in S}(\text{MLP}(s)))$
Note that, although we experimented primarily with the architecture presented above, the framework is general enough to be used with any other hypergraph network. Our experiments demonstrated that even simpler hypergraph models enable the discovery of accurate higher-order structures.
@@ 4 Experimental Analysis
While the importance of higher-order processing is widely accepted in the machine learning community [49] and beyond [50], the amount of benchmarks developed to properly validate the hypergraph methods is still insufficient. This issue becomes even more pronounced when it comes to latent hypergraph inference. Although there is evidence that many existing real-world tasks exhibit underlining higher-order interactions that can be beneficial to capture, evaluating the capability of a neural network to predict the latent structure remains challenging. First, the input features should contain enough information to predict the hypergraph structure. Secondly, even if we do not need the hypergraph structure as a supervision signal, having access to it is necessary to properly evaluate to what extent the model learns to infer the correct structure.
@@
In our work we adopt both of these approaches. First, we will perform an in depth ablation study on a synthetic dataset containing particle simulations. Then, we will move on to the NBA real-world dataset where we will evaluate to what extent our latent hypergraph structure improves the trajectory prediction performance. Our goal is two fold. We show that our model is capable of learning an appropriate latent hypergraph structure in an unsupervised way. Additionally, we prove that the latent higher-order structure, learnt jointly with the rest of the model, helps the performance of the downstream task.
@@ 4.1 Particle Simulations
Figure 2: Ablation studying the importance of hypergraph inference on the Particle Simulation. The sequential prediction and constrained $k$-subset sampling clearly helps the performance. SPHINX obtains the best results across both datasets.
Figure 3: Predicted trajectories and the latent hypergraph structures on the Particle Simulation and the real-world NBA datasets. The model learns to produce accurate trajectories and useful hypergraph structures (depicted as polygons) that correlates well with the ground truth connectivity. We observe that, for the NBA dataset, the hyperedges tends to contain the ball-node (represented as a white dot). This aligns well with our intuition that players dynamics should be highly influence by the position of the ball.
@@
Our simulated system consist of $N$ particles moving in a 2D space. For each example, $K$ random triangles were uniformly sampled to represent $K$ 3-order interactions. All the particles that are part of a triangle rotate around the triangle’s center of mass with a random angular velocity $\theta_{i}$ characteristic to each particle $i$. If one particle is part of multiple triangles, the rotation will happen around their average center of mass. Examples of such trajectories are depicted in Figure 3. For each trajectory we observe the position of the particles in the first $22$ steps, and the task is to infer the trajectory for the following $25$ steps. We are experimenting with 2 variants of the dataset. A simpler one, containing a single higher-order interaction per trajectory (denoted as One-Triangle), and a more challenging one containing two higher-order interaction per trajectory (denoted as Two-Triangles). Each dataset contains $1000$ training, $1000$ validation and $1000$ test trajectories.
@@
For all our experiments, the hypergraph predictor treats particles as nodes. For each particle, the node features consist of the $(x,y)$ coordinates corresponding to the first $T$ timesteps. These features will be used to predict the hypergraph incidence matrix $H$. During training, the hypergraph processor receives the predicted hypergraph structure together with the particle’s position at timestep $t$, and the goal is to predict the position of the particles at timestep $t+1$. To generate longer trajectories, during testing, the model receives at each moment in time the position predicted at the previous timestep.
@@
Hypergraph discovery. These datasets allow us to experiment with learning the higher-order relations in a setup where we have access to the ground-truth connectivity. Our model produces discrete hypergraph structures that we can inspect and evaluate, offering an additional level of interpetability to the framework. Visualisations of our learned hypergraph structure (see Figure 3) reveals that they are highly correlated to the ground-truth interactions used to generate the dataset.
@@
We quantitatively evaluate to what extent our model learns the true higher-order relationships, by computing the overlap between our model’s prediction and the ground-truth hyperedges. We refer to the Supplementary Material for a full descriptions of the metric used in this experiments. In Figure 4 we observe that, even if we do not explicitly optimise for this task, the accuracy of the hypergraph structure predictor increases during training, reaching more than $90\%$ overlap for both the One-Triangle dataset and the Two-Triangles dataset.
@@
General and versatile, our model is designed to predict hypergraph structure that is useful for any hypergraph neural network. However, since the supervision signal comes from the high-level tasks, the architecture used for hypergraph processing can impose certain inductive biases, influencing the hypergraph structure learned by the predictor. In the experiment depicted in Figure 4, we investigate to what extent our model learns an accurate higher-order structure, regardless of the hypergraph processing architecture. We train a set of models using the same hypergraph predictor for inferring the hypergraph structure, but various hypergraph networks for processing it. For this, we experiment with AllDeepSets [28], HGNN [26] and HCHA [31]. The results show that our model is capable of learning the suitable structure irrespective of the specifics in the processing model.
These experiments prove that the hypergraph discovery task is highly correlated with the trajectory prediction problem. Our model, explicitly designed to infer the latent hypergraph structure, is capable not only to accurately solve the downstream task but also to recover the true higher-order relations, without access to the ground-truth hypergraph connectivity.
@@
Importance of sequential prediction. While the classical slot-attention algorithm infers all the clusters in parallel, we argue that this is not an appropriate design choice for hypergraph prediction. Often, the set of hyperedges is mostly symmetric, which makes the problem of attaching the slots to different hyperedges highly ambiguous. In fact, we visually observed that, when predicting all the hyperedges simultaneously, the model tends to associate multiple slots to the same hyperedge, leaving others completely uncovered (see Supplementary Material for more visualisations). This behaviour leads to both a decrease in downstream performance and a less accurate hypergraph predictor. As described in Section 3.1, SPHINX alleviates this issue by inferring the slots sequentially, and providing at each step historical information about the structure predicted so far. Figure 3 and Table 1 demonstrates the benefits of our design choices. Sequential prediction have a clear advantage, outperforming the classical slot-attention predictor.
@@
Influence of $k$-sampling. Previous work on inferring hypergraphs [43] takes inspiration from the graph relational inference [19] where the differentiable sampling is achieved using the Gumbel-Softmax trick. We argue that, inferring the subset of nodes that are part of a hyperedge generates more challenges compare to the graph scenarios. In the beginning of training, sampling each node-hyperedge incidence independently, without any constraints, can lead to hyperedges containing either too few or too many nodes, which highly damage the optimisation. In this work, we leverage the recent advance in differentiable $k$-subset sampling, which imposes the $k$-nodes constraint, while still allowing to estimate the gradient during backpropagation. The experiments in Figure 3 show that the $k$-subset sampling improves the results, leading to an easier optimisation and better final performance. This is in line with the results reported in [43] where smoothness and sparsity regularisation are crucial for improving the results on a real-world dataset. In contrast, our model obtains competitive performance without any optimisation trick.
Comparison with baselines. Both versions of the datasets used in our experiments heavily relies on higher-order processing. In order to predict the particle movement, the model needs to identify the higher-order structures that guides the rotation and use it to generate the displacement from the previous position. To understand to what extent our model is able to produce and process a useful hypergraph structure, we compare against two baseline: a model using a random structure (denoted as random in Figure 3) and a node-level model, that ignores the hypergraph structure (denoted as w/o structure) by setting the incidence matrix on zero. Both models perform much worse that our learnable hypergraph model, indicating that the dataset highly depends on higher-order interactions and our model is capable of discovering an appropriate latent structure. Inspired by [19], we also compare our model against an LSTM baseline that is agnostic to the hypergraph structure and only receives the concatenation of all the particles. While performing comparable on the 1 step prediction metric (2e-5 MSE for SPHINX and 3e-5 for the LSTM), the performance degrades quickly when evaluating on the long-range prediction (6e-3 MSE for SPHINX compared to 1.7e-1 for the LSTM). For the complete numerical results, see Supplementary Material.
Figure 4: Hypergraph discovery improves during training on the synthetic datasets, even if the model is not supervised for this task. Regardless of the hypergraph architectures, the overlap between our (unsupervised) predicted hypergraph and the gt. connectivity increases, demonstrating that the model recovers the real structure to a high extent.
@@
Table 1: Ablation study on the NBA dataset. Both using sequential, history-aware predictors and using $k$-subset sampling prove to be beneficial for the final performance (ADE/FDE metrics reported). Moreover, our model shows a clear advantage compared to the best recent work on the single-trajectory prediction task.
Time GroupNet SPHINX SPHINX SPHINX
w/o sequential w Gumbel
1sec 0.65/1.03 0.62/0.94 1.29/1.81 0.59/0.92
2sec 1.38/2.61 1.18/2.08 2.10/3.34 1.12/2.06
3sec 2.15/4.11 1.73/3.14 2.81/4.60 1.65/3.13
4sec 2.83/5.15 2.25/4.08 3.45/5.66 2.14/4.09
@@
Implementation details. Unless otherwise stated, the One-Triangle experiments use a single slot in the slot-attention module, while the Two-Triangles uses two slots. For estimating the gradients in the $k$-subset sampling algorithm we experiment with SIMPLE [8], AIMLE [9] and IMLE [46] algorithms. According to our experiments, the first two perform on par, while IMLE [46] mostly suffer from gradient issues. Unless otherwise specified, our models use the AllDeepSets [28] hypergraph processor. In all experiments, we use Adam optimizer for 1000 epochs, trained on a single GPU.
@@
4.2 NBA Dataset

### v2

@@ 3.2 Discrete constrained sampling
The hypergraph predictor module, as described above, produces a probabilistic incidence matrix $\mathbf{P}\in\mathbb{R}^{N\times M}$, where each element $p_{i,j}$ denotes the probability of a node $i$ being part of the hyperedge $j$. However, the standard hypergraph neural network architectures are designed to work with sparse and discrete rather than probabilistic structures.
Previous work (Li et al., 2022) employs Gumbel-Softmax (Jang et al., 2017) to differentiably sample from a categorical distribution. However, these techniques sample each element in the hyperedge independently, without any control on the cardinality of the hyperedge. This leads to unstable optimisation, that requires additional training strategies such as specific sparsity regularisation. Top-k is another popular approach to achieve sparsity (Zhou et al., 2023; Xu et al., 2022). However, to preserve gradients, the generated structure cannot be discretized. Although a weighted hypergraph can be beneficial for certain architectures, it is not universally compatible with all decoders and makes the interpretation of the latent space more challenging.
To address this issue, we are leveraging the recent advancement in constrained $k$-subset sampling (Ahmed et al., 2023; Niepert et al., 2021; Minervini et al., 2023). These methods were successfully used to tackle discrete problems such as combinatorial optimisation, learning explanations and, more recently, rewiring graph topology (Qian et al., 2023). Different from classical differentiable samplers (Jang et al., 2017), the $k$-subset sampler would produce a subset of size exactly $k$, equipped with a gradient estimator useful for backpropagation.
@@
Figure 2: Ablation studying the importance of hypergraph inference on the Particle Simulation. (a) The sequential prediction and constrained $k$-subset sampling clearly helps the downstream performance. (b) SPHINX also predicts more accurate hypergraphs in terms of overlap. (c) Moreover, hypergraph discovery improves during training on the synthetic datasets, even if the model is not supervised for this task. Regardless of the hypergraph architectures, the models achieve a high overlap between the (unsupervised) predicted hypergraph and the gt. connectivitys.
@@ 3.3 Hypergraph processing
Producing a discrete hypergraph structure and being able to propagate the gradient through the entire pipeline, enable us to process the resulting latent hypergraph with any existing architecture designed for higher-order representations. In recent years, several architectures were developed for hypergraph-structured input (Feng et al., 2019; Wang et al., 2022; Huang & Yang, 2021; Chien et al., 2022). Most of them follow the general two-stage message-passing framework. In the first stage, the information is sent from nodes to the hyperedges using a permutation-invariant operator $z_{j}=f_{V\rightarrow E}(\{x_{i}|v_{i}\in e_{j}\})$. In the second stage the messages are sent back from hyperedge to nodes $x_{i}=f_{E\rightarrow V}(\{z_{j}|v_{i}\in e_{j}\})$.
In our experiments we use a similar setup to the one proposed in (Chien et al., 2022), in which the two functions $f_{V\rightarrow E}$ and $f_{E\rightarrow V}$ are implemented as DeepSets (Zaheer et al., 2017) models $f_{V\rightarrow E}(S)=f_{E\rightarrow V}(S)=\text{MLP}(\sum_{s\in S}(\text{MLP}(s)))$. Note that, although we experimented primarily with the architecture presented above, the framework is general enough to be used with any other hypergraph network. Our experiments demonstrated that even simpler hypergraph models enable the discovery of accurate higher-order structures.
@@ 4 Experimental Analysis
While the importance of higher-order processing is widely accepted in the machine learning community (Zhou et al., 2006) and beyond (Zhang et al., 2020), the amount of benchmarks developed to properly validate the hypergraph methods is still insufficient. This issue becomes even more pronounced when it comes to latent hypergraph inference. Although there is evidence that many existing real-world tasks exhibit underlining higher-order interactions that can be beneficial to capture, evaluating the capability of a neural network to predict the latent structure remains challenging. First, the input features should contain enough information to predict the hypergraph structure. Secondly, even if we do not need the hypergraph structure as a supervision signal, having access to it is necessary to properly evaluate to what extent the model learns to infer the correct structure.
@@
In our work we adopt both approaches. First, we perform an in-depth ablation study on a synthetic dataset containing particle simulations. Then, we move to the real-world datasets where we evaluate on both inductive and transductive setups. Our goal is two-fold. We show that our model is capable of learning an appropriate latent hypergraph structure in an unsupervised way. Additionally, we prove that the latent higher-order structure, learnt jointly with the rest of the model, helps the performance of the downstream task on both the inductive and transductive problems.
@@ 4.1 Particle Simulations
Our simulated system consists of $N$ particles moving in a 2D space. For each example, $K$ random triangles were uniformly sampled to represent $K$ 3-order interactions. All the particles that are part of a triangle rotate around the triangle’s center of mass with a random angular velocity $\theta_{i}$ characteristic to each particle $i$. If one particle is part of multiple triangles, the rotation will happen around its average center of mass. Examples of such trajectories are depicted in Figure 3 and Figure 6. For each trajectory we observe the position of the particles in the first $22$ steps, and the task is to infer the trajectory for the following $25$ steps. Each example exhibits two higher-order interactions. The dataset contains $1000$ trajectories for training, $1000$ for validation and $1000$ for test.
@@
For all experiments, the hypergraph predictor treats particles as nodes. For each particle, the node features consist of the $(x,y)$ coordinates corresponding to the first $T$ timesteps. These features will be used to predict the incidence matrix $H$. During training, the hypergraph processor receives the predicted hypergraph structure together with the particle’s position at timestep $t$, and the goal is to predict the position of the particles at timestep $t+1$. To generate longer trajectories, during testing, the model receives at each moment in time the position predicted at the previous timestep.
@@
Hypergraph discovery. These datasets allow us to experiment with learning the higher-order relations in a setup where we have access to the ground-truth connectivity. Our model produces discrete hypergraph structures that we can inspect and evaluate, offering an additional level of interpretability to the framework. Visualisations of our learned hypergraph structure (see Figure 3) reveal that they are highly correlated to the ground-truth interactions used to generate the dataset.
@@
We quantitatively evaluate to what extent SPHINX learns the true higher-order relationships, by computing the overlap between our model’s prediction and the ground-truth hyperedges. We refer to the Appendix for a full description of the metric used in this experiment. In Figure 2.c we observe that, even if we do not explicitly optimise for this task, the accuracy of the hypergraph structure predictor increases during training, reaching up to $90\%$ overlap.
@@
General and versatile, our model is designed to predict hypergraph structure that is useful for any hypergraph neural network. However, since the supervision signal comes from the high-level tasks, the architecture used for hypergraph processing can impose certain inductive biases, influencing the hypergraph structure learned by the predictor. In the experiment depicted in Figure 2.c, we investigate to what extent our model learns an accurate higher-order structure, regardless of the hypergraph processing architecture. We train a set of models using the same hypergraph predictor for inferring the hypergraph structure, but various hypergraph networks for processing it. For this, we experiment with AllDeepSets (Chien et al., 2022), HGNN (Feng et al., 2019) and HCHA (Bai et al., 2021). The results show that our model is capable of learning the suitable structure irrespective of the specifics in the processing model.
@@
Importance of sequential prediction. While the classical slot-attention algorithm infers all the clusters in parallel, we argue that this is not an appropriate design choice for hypergraph prediction. Often, the set of hyperedges is mostly symmetric, which makes the problem of attaching the slots to different hyperedges highly ambiguous. In fact, we visually observed that, when predicting all the hyperedges simultaneously, the model tends to associate multiple slots to the same hyperedge, leaving others completely uncovered (see SPHINX w/o sequential model in Figure 6). This behaviour leads to both a decrease in downstream performance and a less accurate hypergraph predictor. As described in Section 3.1, SPHINX alleviates this issue by inferring the slots sequentially, and providing at each step historical information about the structure predicted so far. Figure 2.a and 2.b demonstrate the benefits of our design choices. Sequential prediction (SPHINX) has a clear advantage, outperforming the classical simultaneous predictor (SPHINX w/o sequential).
@@
Influence of $k$-sampling. Previous work on inferring hypergraphs (Li et al., 2022) takes inspiration from the graph domain (Fetaya et al., 2018) where the differentiable sampling is achieved using the Gumbel-Softmax trick. We argue that, inferring the subset of nodes that are part of a hyperedge generates more challenges compares to the graph scenarios. At the beginning of training, sampling each node-hyperedge incidence independently, without constraints, can lead to hyperedges containing either too few or too many nodes, which highly damages the optimisation. We can observe this phenomenon visually in Figure 6, column SPHINX w gumbel. In this work, we leverage the recent advance in differentiable $k$-subset sampling, which imposes the $k$-nodes constraint, while still allowing us to estimate the gradient during backpropagation. The experiments in Figure 2.a and Figure 2.b (see gumbel model vs SPHINX model) show that the $k$-subset sampling improves the results, leading to an easier optimisation and better final performance. This is in line with the results from (Li et al., 2022) where smoothness and sparsity regularisation are crucial for improving the results on a real-world dataset. In contrast, our model obtains competitive performance without any optimisation trick.
@@
Table 1: Ablation study on the NBA dataset. Both using sequential, history-aware predictors and using $k$-subset sampling prove to be beneficial for the final performance. Moreover, SPHINX shows a clear advantage compared to the existing inductive hypergraph predictors on the single-trajectory prediction task (ADE/FDE metrics).
1sec 2sec 3sec 4sec
SPHINX w/o sequential 0.62/0.94 1.18/2.08 1.73/3.14 2.25/4.08
SPHINX w Gumbel 1.29/1.81 2.10/3.34 2.81/4.60 3.45/5.66
TDHNN 0.68/1.03 1.27/2.22 1.84/3.31 2.30/4.29
GroupNet 0.65/1.03 1.38/2.61 2.15/4.11 2.83/5.15
SPHINX 0.59/0.92 1.12/2.06 1.65/3.13 2.14/4.09
Comparison with inductive hypergraph predictors. The Particle Simulation dataset heavily relies on higher-order processing. In order to predict the particle movement, the model needs to identify the higher-order structures that guide the rotation. To understand to what extent our model is able to produce and process a useful hypergraph structure, we compare against two baselines: a model using a random structure (denoted as random in Figure 2) and a node-level model, that ignores the hypergraph structure (denoted as w/o structure). Both models perform much worse than our learnable hypergraph model, indicating that the dataset highly depends on higher-order interactions and that our model is capable of discovering an appropriate latent structure.
We also compared against existing hypergraph predictors in the literature. Most of the existing methods are constrained to only work in the transductive setup (by only predicting a single hypergraph for all examples). The only models that are applicable in the inductive setup are GroupNet (Xu et al., 2022), TDHNN (Zhou et al., 2023) and EvolveHypergraph (Li et al., 2022) [^2] EvolveHypergraph does not have public code.. The results in Figure 2.a show that our predictor outperforms all the existing ones. We believe that the drop in performance for the existing models is due to the poor gradients used in updating the predictor. Both GroupNet and TDHNN models use a form of top-k selection which only partially propagates the gradients. Moreover, qualitative evaluation shows that TDHNN suffers from the same limitation as our "w/o sequential baseline": all the hyperedges ended up selecting the same set of nodes.
Our model not only outperforms the existing ones in terms of downstream performance, but also discovers more accurate hyperedges. Computing the overlap between the prediction and the oracle hyperedges, SPHINX achieves an overlap of $90\%$ compared to $67\%$ for TDHNN (Figure 2.b) [^3] For GroupNet we can not compute the overlap since the number of edges is constrained to be equal to the number of nodes..
Table 2: Performance on the transductive datasets using various features. SPHINX obtains better results compared to both the static methods, where the structure doesn’t modify during training, and dynamic methods, where the hypergraph is optimised. Mean/std computed over $20$ runs.
NTU ModelNet40
Model GVCNN MVCNN GV+MV GVCNN MVCNN GV+MV
Static GCN 78.80$\pm$0.92 78.72$\pm$1.97 80.43$\pm$1.09 91.80$\pm$0.46 91.50$\pm$1.80 94.85$\pm$1.75
GAT 79.60$\pm$0.03 78.50$\pm$1.17 80.16$\pm$1.08 91.65$\pm$0.25 90.07$\pm$0.41 95.75$\pm$0.14
HGNN 82.50$\pm$1.62 79.10$\pm$0.90 83.64$\pm$0.37 91.80$\pm$1.73 91.00$\pm$0.66 96.96$\pm$1.43
HGNN+ 82.80$\pm$1.11 76.40$\pm$1.17 84.18$\pm$0.82 92.50$\pm$0.08 90.60$\pm$1.68 96.92$\pm$1.81
HNHN 83.10$\pm$1.89 79.60$\pm$0.79 80.60$\pm$0.95 92.10$\pm$1.76 91.10$\pm$1.84 93.80$\pm$1.84
HyperGCN 79.90$\pm$1.78 78.10$\pm$0.83 79.90$\pm$0.91 92.20$\pm$0.80 90.20$\pm$0.28 96.10$\pm$0.63
Dynamic DHGNN 82.30$\pm$0.98 77.60$\pm$1.55 85.13$\pm$0.26 92.13$\pm$1.55 85.53$\pm$0.83 96.99$\pm$1.46
DeepHGSL 76.28$\pm$1.45 72.30$\pm$1.57 78.67$\pm$0.77 89.32$\pm$0.71 88.62$\pm$0.93 90.33$\pm$0.66
HSL 81.82$\pm$1.30 75.68$\pm$1.41 82.26$\pm$1.20 93.17$\pm$0.25 91.44$\pm$0.42 96.92$\pm$0.41
TDHNN* 92.70$\pm$1.61 90.76$\pm$1.05 92.70$\pm$1.26 96.96$\pm$0.34 91.92$\pm$0.50 98.69$\pm$0.30
SPHINX 94.80$\pm$0.90 92.95$\pm$0.85 94.67$\pm$0.71 97.29$\pm$0.29 92.79$\pm$0.41 98.92$\pm$0.21
Figure 3: Predicted trajectories and the latent hypergraph on two examples of the Particle Simulation.
@@
Implementation details. Unless otherwise stated, experiments use two slots and the AllDeepSets model as a decoder. For estimating the gradients in the $k$-subset sampling algorithm we experiment with SIMPLE (Ahmed et al., 2023), AIMLE (Minervini et al., 2023) and IMLE (Niepert et al., 2021) algorithms. The first two perform on par, while IMLE suffers from gradient issues. We use Adam optimizer for 1000 epochs, trained on a single GPU.
@@
Table 3: Performance in the inductive setup. in terms of $\text{minADE}_{20}$ and $\text{minFDE}_{20}$ metrics for the NBA dataset.
Method 1sec 2sec 3sec 4sec
Graph NRI 0.51/0.74 0.96/1.65 1.42/2.50 1.86/3.26
DNRI 0.59/0.70 0.93/1.52 1.38/2.21 1.78/2.81
EvolveGraph 0.35/0.48 0.66/0.97 1.15/1.86 1.64/2.64
STGAT 0.45/0.66 0.87/1.41 1.28/2.08 1.69/2.66
Trajectron++ 0.44/0.67 0.79/1.18 1.51/2.49 2.09/3.52
HyperG. EvolveH (H) 0.49/0.74 0.95/1.68 1.44/2.27 1.91/3.08
EvolveH (H+G) 0.33/0.49 0.63/0.95 0.93/1.36 1.21/1.74
TDHNN 0.68/1.03 1.27/2.22 1.84/3.31 2.30/4.29
GroupNet 0.34/0.48 0.62/0.95 0.87/1.31 1.13/1.69
SPHINX 0.30/0.43 0.59/0.94 0.88/1.38 1.16/1.74
4.2 Inductive setup